# QADM-Net: Multi-Level Quality-Adaptive Dynamic Network for Reliable Multimodal Classification

### v2

@@ 4.2.4 Informative features that QADM-Net learns (Q4)
We further demonstrate the reliability of QADM-Net by analyzing the informative features it retains following previous works Han et al. (2022); Zou et al. (2023). Table 4 shows the top 5 informative biomarkers of the three modalities on the BRCA dataset retained by the feature informativeness masks $q_{i}^{m,t}$ (Equation 2) in QADM-Net, which align with the findings of existing biomedical research. For example, Low expression of ZNF671 Zhang et al. (2019), miR-378 Arabkari et al. (2023) are both associated with poor prognosis in breast cancer. KLK8 is an independent indicator of prognosis in breast cancer patients Michaelidou et al. (2015).
4.2.5 Computation Effectiveness of QADM-Net (Q5)
@@
Although computational efficiency is not the design motivation or novelty of QADM-Net, it maintains the high computational efficiency characteristic of dynamic networks while ensuring model robustness. Table 5 compares the classification accuracy and corresponding FLOPs of QADM-Net with some SOTA reliable multimodal classification methods and the multimodal dynamic neural network DynMM on both clean and noisy datasets. QADM-Net achieves most stable and promising results with only a 5.12M increase in FLOPs (less than DynMM and still lower than GCFANet after the increase), demonstrating its computation efficiency while ensuring model reliability. Futher comparison of the parameter amounts can be found in the supplementary material.
5 Conclusions
@@
This work proposes a novel multi-level quality-adaptive dynamic multimodal network (QADM-Net) for reliable multimodal classification. On the one hand, QADM-Net adopts a reliable multi-level quality estimation method NFCE based on noise-free prototypes and classifier-free design. On the other hand, QADM-Net proposed two mechanisms termed GCND and LGP to achieve sample-adaptive network depth and parameters based on modality and feature level quality, significantly promoting inference reliability. Extensive experiments show the superiority of QADM-Net in classification performance compared to state-of-the-art reliable classification methods and demonstrate the strong adaptability of QADM-Net to data with diverse quality.

### v3

@@ IV-G Informative features that Multi-QuAD learns
The reliability of Multi-QuAD is further demonstrated by analyzing the informative features it retains following previous works [4, 28]. Table VI shows the top 5 informative biomarkers of the three modalities on the BRCA dataset retained by the feature informativeness masks $q_{i}^{m,t}$ learned in Eq.(2) by Multi-QuAD, which align with the findings of existing biomedical research. For example, low expression of ZNF671 [64], miR-378 [65] are both associated with poor prognosis in breast cancer. KLK8 is an independent indicator of prognosis in breast cancer patients [66]. Furthermore, t-SNE visualizations on the features from different modalities of all samples $\{\{f_{i}^{m}\}_{m=1}^{M}\}_{i=1}^{N}$ are performed to show the informativeness of the features that Multi-QuAD learns. Fig. 11a and 11b show the results on the BRCA and ROSMAP datasets. The features of all modalities enhanced by Multi-QuAD exhibit minimal intra-class distances and clear inter-class separation, which indicate that Multi-QuAD effectively retains informative features and thus achieves reliability.
TABLE VI: Top 5 informative biomarkers that Multi-QuAD learned from different modalities on BRCA dataset.
Modality Top 5 informative biomarkers
DNA ZNF671, KRTAP3-1, ZCCHC8, AGR2, MYT1
mRNA KLK8, MIA, PTX3, SOX11, KRT6B
miRNA mir-378, mir-190b, mir-934, mir-21, mir-23b
TABLE VII: Comparison of ACC, FLOPs per sample, and #param of different methods on the FOOD101 dataset, where * are the results under Gaussian noise with $\sigma=5$.
Method GCFANet PDF DynMM Multi-QuAD ACC (%) 92.9 93.3 89.3 94.5 FLOPs (M) 9.22 3.53 3.71 3.82 #param (M) 9.13 3.46 3.67 3.77 ACC (%) 89.4 (-3.5) 90.1 (-3.2) 70.1 (-19.2) 93.8 (-0.7) FLOPs (M) 9.22 (+0) 3.53 (+0) 10.82 (+7.11) 8.94 (+5.12) #param (M) 9.13 (+0) 3.46 (+0) 10.53 (+6.86) 8.91 (+5.14)
IV-H Computation Effectiveness of Multi-QuAD
@@
Extensive experiments also demonstrate that Multi-QuAD achieves high computational efficiency while ensuring reliability. Table VII compares the classification accuracy (ACC), FLOPs, and parameter amount (#param) of Multi-QuAD with some SOTA reliable multimodal classification methods and the multimodal dynamic neural network DynMM on both clean and noisy datasets. Multi-QuAD achieves the most stable and promising results with only a 5.12M increase in FLOPs and a 5.14M increase in #param (less than DynMM and still lower than GCFANet after the increase) under noise.
V Conclusions
@@
This work proposes a novel multi-level quality-adaptive dynamic multimodal network (Multi-QuAD) for reliable multimodal classification. Multi-QuAD adopts a reliable multi-level quality estimation method, NFCE, based on noise-free prototypes and classifier-free design. Additionally, Multi-QuAD proposes two mechanisms, termed GCND and LGP, to achieve sample-adaptive network depth and parameters based on modality and feature-level quality, thereby significantly enhancing inference reliability. Extensive experiments on four datasets have demonstrated Multi-QuAD’s significant superiority over other state-of-the-art methods in terms of classification performance and reliability. The adaptability of Multi-QuAD to diverse data quality, the effectiveness of its components (NFCE, GCND, and LGP), and its high computational efficiency have also been verified and discussed.
Existing multimodal learning methods have not been sufficiently studied for their reliability in other machine learning tasks, such as regression and semantic segmentation. This represents a novel and worthwhile problem for in-depth exploration. Therefore, future research in this article will be further extended and validated on a more diverse range of machine learning tasks.